amortization table

Saturn issues 6.5%, five-year bonds dated January 1, 2011, with a $500,000 par value. The bonds pay interest on June 30 and December 31 and are issued at a price of$510,666. The annual market rate is 6% on the issue date.

2. Prepare an effective interest amortization table for the bonds’ life.

The checked answer in the book is 6/30/2013 for Carrying Value is $505,728.
Answer #1
Concepts and reason

Bond: Bond is a long-term financial instrument issued by a company for which the company pays interest to the bond holders and repays the borrowed amount after a specific period of time.

Amortisation: Amortisation is a process of allocating intangible asset cost throughout the useful life or the period which the asset exists. It also refers to payment of debt for certain time inclusive of instalments.

Transaction: A transaction in business refers to any events that affect the financial position of the business that can be reliably measured in monetary terms. A business transaction affects the accounting equation.

Journal entry: The record of business transactions in a chronological order with equal amounts of debits and credits in the journal is referred to as journal entry.

General Journal: It is a two-column journal book in which the financial transactions that takes place in the company’s day-to-day business activities are recorded in a chronological order.

Fundamentals

Interest expense: This is a non-operating expense and it will be shown on the income statement. Generally, this expense is considered as an amount incurred by the company for using the other resources like borrowing loan from bank or purchasing a note etc.

Face value: Face value is the original price of a stock or bond. It is also called as par value.

Issue price of the bond: The sum of present value of interest payments (annuity) and principal (single sum) is referred to as issue price of bond. It is otherwise known as selling price of the bond.

Bond sells at premium: It means that the bond is issued at a higher price than the face value of the bond. As well as, when the market rate of interest is more than the coupon rate is called as bond issued at premium.

Bond sells at par: When bonds are issued at the actual value neither by discount nor by premium is said to be as bond sold at par.

Rules for debit and credit:

When asset increases debit it; asset decreases credit it.

When liabilities increase credit it; liabilities decreases debit it.

When stockholders’ equity increases credit it; stockholders’ equity decreases debit it.

When expenses and losses increase debit it; expenses and losses decreases credit it.

When incomes and gains increase credit it; incomes and gains decreases debit it.

1)

Compute the total bond interest expense over the bonds life:

Totalinterestexpenses}=(Facevalue×Rate×PeriodPremium)=($250,000×6.5%×612×(5×2)$5,333)=$81,250$5,333=$75,917.\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total interest }}\\\\{\rm{expenses }}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Face value }} \times {\rm{ Rate }} \times {\rm{ }}\\\\{\rm{Period }} - {\rm{ Premium}}\\\end{array} \right)\\\\{\rm{ = }}\left( \begin{array}{l}\\{\rm{\$ 250,000 }} \times {\rm{ 6}}{\rm{.5\% }} \times \frac{6}{{12}}{\rm{ }} \times {\rm{ }}\\\\\left( {{\rm{5 }} \times {\rm{ 2}}} \right){\rm{ }} - {\rm{ \$ 5,333}}\\\end{array} \right)\\\\ = \,{\rm{ \$ 81,250 }} - {\rm{ \$ 5,333}}\\\\{\rm{ = \$ 75,917}}{\rm{.}}\\\end{array}

Therefore, the total bond interest expenses over the bond’s life are $75,917.

Working notes:

Calculation of premium on bonds is given below:

Premiumonbonds}=IssuepriceFacevalue=$255,333$250,000=$5,333.\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Premium on }}\\\\{\rm{bonds}}\\\end{array} \right\}{\rm{ = Issue price }} - {\rm{ Face value}}\\\\{\rm{ = \$ 255,333 }} - {\rm{ \$ 250,000}}\\\\{\rm{ = \$ 5,333}}{\rm{.}}\\\end{array}

Therefore, the premium on bonds is $5,333.

2)

Prepare the effective interest amortization table:

Semi annual Cash Bond interest Premium Unamortized Carrying
period interest paid expense amortization premium | value
1/1/201

Therefore, the amortization table for bonds life is prepared.

3)

Prepare the journal entry to record the first two interest payments:

Credit
Debit
$7,660
$465
$8,1251
Date Account Title and Explanation
6/30/2015 Interest Expense
Premium on Bonds payable
Cash

Therefore, the journal entry is passed for the first two interest payments.

Ans: Part 1

The total bond interest expenses over the bond’s life are $75,917.

Part 2

Semi annual Cash Bond interest | Premium Unamortized Carrying
period interest paid expense amortization premium | value
1/1/2

Part 3

Credit
Debit
$7,660
S465
$8,125
Date Account Title and Explanation
6/30/2015 Interest Expense
Premium on Bonds payable
Cash
(